July 13 OWL shift summary

Richard Kessler (rkessler@fnal.gov)
Tue, 13 Jul 1999 06:57:33 -0500 (CDT)

Shift Summary for OWL shift, July 13, 1999

On Shift:RK SET

Hours of beam delivered at all intensities:
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:
- 15 minutes for Cr 53 full and repeater power cycle,
and ocassional ktevana errors
- 28 seconds of DAQ dead time
( > scaler_screen 13870 daqdead_sec)

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Easy data taking shift. Beam is very consistent.

Discovered that the RC10 trigger source is toast since run 13861 but
was not caught (see trg#veto or #summary macro). The "extra" RC trigger
sources are fine so offline analysis will be OK; hence no need to panic.
RC10 source will get fixed during tuesday shutdown.

Also discovered the KQ4Y chan 43 is OK in run 13870.
On the other hand DC3Y has weak wires 81 & 82 since run 13867.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13870 running smoothly at start of shift.

182 ktevana errors while investigatng RC10 problem
at RR160 (and unable to hear audio alarms);
probably lost around 4 spills.

205 Cr 53 FULL; kfrend to the rescue. Lost 2 spills

218-221 4:35am ktevana errors. Kfrend works only after cycling
the downstairs repeater. Lost four spills.

5:45 MCR calls to warn that beam goes off at 6:15am.
6:30 beam goes away. STOP run

13871 6:45 Begin LASER_SCAN

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift:

- PAW > trg#veto (to watch RC10 trigger source)